# Mamba-based Deep Learning Approaches for Sleep Staging on a Wireless Multimodal Wearable System without Electroencephalography

### v2

@@ 4.3 Model Performance
Both CRNN and RNN model architectures obtain maximal per-class performances for wake and REM, and minimal per-class performances for N1 and N2 sleep. As a transitory state, periods of N1 sleep are short and instable, making its discrimination a difficult task. Physiological similarities between N2 and N3 sleep, such as cardiorespiratory signatures, similarly complicate the model’s ability to differentiate them, as evidenced by N2-N3 confusion as the most frequent misprediction for 5-class models (as well as N1/N2-N3 confusion for 4-class models). Ultimately, the differences betwen N1, N2, and N3 represent a matter of depth rather than kind, and some degree of difficulty in distinguishing these without EEG is to be expected. Conversely, wake and REM sleep have clearer differentiating characteristics, such as higher heart rate, which allows for easier classification.
In spite of a simpler feature set and smaller parameter count, the RNNs outperform the CRNNs in virtually every metric. It is possible that the independent processing of segmented windows by the CNN component interfered with the wider CRNN’s ability to discern changes at neighbouring window boundaries. As well, during epochs where sleep is transitioning between classes, the provided label becomes less reliable; the CRNN attempts to predict a behaviourally inconsistent epoch while it is likely the RNN more flexibly captures this ambiguity by means of multiple differing fine-resolution class predictions throughout the 30 seconds.
No statistically significant variations in accuracy based on age, sex, PLMI, BMI, TST, or sleep efficiency were exhibited by the RNN 5-class ensemble model, providing an important advantage when applied in clinical contexts. The marginally significant () association between AHI and accuracy may owe to the fact that individuals with sleep apnea experience sleep fragmentation, resulting in N1 sleep, one of the most difficult stages to separate; nonetheless, the association was modest ().
The series of ablation analyses establish that much of the RNN model’s performance was likely driven by chest-module features (i.e. ECG, temperature, and xyz accelerometry) rather than limb module (i.e. PPG and temperature), meaning that omitting the limb module may be reasonable if participant burden is a concern and collection of oximetry data is not necessary. Remarkably, exclusion of both ECG and PPG data with retention of only accelerometry and temperature data resulted in model performance almost indistinguishable from use of all the chest module sensors, suggesting that the inclusion of traditional heart rate sensors (PPG, ECG) is not necessary. This highlights the richness of accelerometry data obtained at the chest, spectrograms of which may capture both cardiac (ballistocardiogram) and respiratory data, as well as measures of movement. It may also contribute to the robustness of our model given the relative reliability of the accelerometers compared to the ECG/PPG sensors.
@@ 4.4 Limitations, Strengths, and Future Directions
This study had many strengths: 1) The ANNE One system itself, with multiple channels of data recorded from a pair or minimally intrusive sensors, is a strength insofar as it provides an element of redundancy should one sensor return low quality signal, 2) the inclusion of Mamba allowed the achievement of relatively high performance with a relatively parameter-scarce architecture, and 3) the study population was varied, including participants from across a spectrum of sleep disorders.
A few methodological limitations are also worth considering. 1) The differing nature of the CRNN and RNN in preprocessing and predicting (notably, the CRNN’s 30-second window as opposed to the RNN’s 1-second window) make the models difficult to compare; our reconciling of the RNN to make predictions for 30-second windows may not reflect its full performance. 2) The black box nature of CRNNs and RNNs make it difficult to ascertain the specific mechanisms used to infer sleep stage from the sensor data. Further analyses targeted toward interpreting the neural networks may be explored. 3) The recordings were overnight with a brief period of wakefulness before lights out, resulting in a preponderance of sleep vs. wake. It is difficult to know for sure how the model will perform in detecting daytime naps during long ambulatory recordings and further studies with ambulatory recordings in clinical populations will be informative. 4) The recordings were all obtained from a single site. Additional work is needed to assess the extent to which these results will be generalizable to ANNE recordings made in different settings on different patient populations.
@@ 5 Conclusion
This work introduces CRNN and RNN models to automatically perform 3-class, 4-class, and 5-class sleep staging using data from the ANNE One dual sensor using sleep stages inferred from concurrent PSG recordings, with the RNN ensemble model topping the others in virtually every performance metric. The model meets or surpasses metrics in related literature, especially for wearable devices, and demonstrates robustness across a wide range of clinical characteristics. Moreover, the model’s capability was achieved without prescreening and excluding large numbers of patients or epochs for reasons of signal quality. Compared to PSGs, the ANNE One sensor is far less intrusive of a device and procedure, which is significant for expanding the accessibility of sleep measurement for older and remote-living adults at greater risk of sleep-related illness. Our model provides a scalable, accurate minimally-intrusive approach to the ambulatory assessment of sleep staging without the need for EEG.
Acknowledgments
@@
Special thanks to Prof. Jimmy Ba from the Department of Computer Science; Prof. Paul Kushner and Prof. Dylan Jones from the Department of Physics at the University of Toronto for offering various forms of support for this research during its infancy.
Appendix A: Summary Statistics by Groups
@@
Appendix B: Bidirectional Mamba Block
Appendix C: Explorations in Feature Selection
A previous iteration of the model used a pure CRNN (GRU) approach, and the approach encompassed two additional strategies, which are utilized within the paper without extensive justification.
The engineered features are a suite of additional features across the time, frequency, and scalar domains used alongside the raw signals as extracted from the ANNE sensors.
Time Domain Frequency Domain Scalar Domain PPG, ECG Heart Rates PPG Freq. Spectra PPG Freq. Spectra Skew PPG, ECG Signal Quality Index ECG Freq. Spectra PPG Time Entropy PPG, ECG Amplitude x-accel. Freq. Spectra PPG Freq. Spectra Entropy Pulse Arrival Time (PAT) y-accel. Freq. Spectra | Heart Rate| Accel. ENMO z-accel. Freq. Spectra Accel. ENMO Kurtosis SpO2 Accel. z-angle Freq. Spectra Entropy Metadata (Gender, Age)
In addition, the auxiliary training objective was introduced in the form of the transition state. These states are derived directly from the target labels. With this, the model attempts to predict whether there will be a change in sleep class from this epoch to the next. The rationale is that sleep stages in reality do not always change at the end of 30-second windows, and many epochs may have data that is transitioning between two separate sleep stages, which would be difficult for the model as they may not resemble other members of their labelled class and instead as a mixture of both sleep stages. By training the model to also predict correct transitions, it should assist the model in determining which epochs may be occurring during a transition between stages, and thus make it more robust in general.
The application of the transition state as an auxiliary objective was attempted in the Mamba approaches in the main paper; however, it did not yield a noticeable increase in performance, and thus was not included in the final models.
The results of the approach are detailed in the graph below, with results indicating that the two approaches mentioned positively impacted results.
3-Class Sleep Staging 2-Class Sleep Detection F1 Precision Accuracy F1 Precision Accuracy Wake Time Series Only Baseline 0.681 0.723 0.644 0.740 0.649 0.860 Engineered Features 0.743 0.723 0.764 0.742 0.703 0.786 Engineered Features & Auxiliary Objective 0.727 0.757 0.700 0.781 0.701 0.880 Ensemble 0.732 0.789 0.683 NREM Sleep Time Series Only Baseline 0.534 0.720 0.425 0.793 0.892 0.714 Engineered Features 0.507 0.754 0.381 0.826 0.859 0.826 Engineered Features & Auxiliary Objective 0.619 0.744 0.530 0.834 0.912 0.769 Ensemble 0.702 0.767 0.647 REM Time Series Only Baseline 0.401 0.265 0.827 Engineered Features 0.428 0.287 0.844 Engineered Features & Auxiliary Objective 0.447 0.314 0.778 Ensemble 0.517 0.381 0.679 Macro Time Series Only Baseline 0.539 0.632 0.555 0.767 0.771 0.770 Engineered Features 0.559 0.663 0.581 0.783 0.784 0.792 Engineered Features & Auxiliary Objective 0.600 0.598 0.624 0.807 0.807 0.811 Ensemble 0.651 0.646 0.679
Appendix D: Scalar and Frequency Features
Appendix E: UMAP of CRNN Model Learned Embeddings
Appendix F: Class-Wise Evaluation Metrics
@@
5-Class Sleep Staging 4-Class Sleep Staging 3-Class Sleep Staging Regular Ensemble Regular Ensemble Regular Ensemble Metric (%) CRNN RNN CRNN RNN CRNN RNN CRNN RNN CRNN RNN CRNN RNN Wake Wake Wake Balanced Accuracy 83.75 84.93 86.47 86.19 84.13 86.12 87.93 87.87 78.57 88.14 87.97 88.27 Precision 84.61 88.16 87.64 88.44 84.48 88.57 88.34 88.67 80.48 88.70 88.24 88.82 Recall 84.68 87.28 87.64 88.04 83.38 88.07 88.32 88.70 76.26 88.72 88.12 88.85 F1 84.63 86.93 87.54 87.81 83.53 87.82 88.33 88.66 76.40 88.71 88.16 88.83 Cohen’s 67.92 72.53 73.93 74.41 66.37 74.42 75.74 76.32 53.63 76.47 75.45 72.89 MCC 67.95 73.94 74.11 75.17 67.04 75.32 75.74 76.37 56.43 76.48 75.49 76.74 N1 Sleep N1/N2 Sleep Non-REM Sleep Balanced Accuracy 62.71 65.07 67.99 69.62 69.22 67.11 71.58 73.47 72.35 83.46 81.03 85.42 Precision 85.30 85.70 86.74 87.09 73.32 68.97 75.37 75.61 74.26 83.60 81.90 85.50 Recall 82.64 78.13 81.74 79.31 73.05 69.43 75.06 75.81 72.28 83.45 80.99 85.41 F1 83.84 81.14 83.77 82.25 71.73 69.01 74.00 75.36 71.73 83.43 80.87 85.41 Cohen’s 21.74 20.73 27.21 26.46 40.67 34.97 45.39 48.25 44.64 66.90 62.01 70.83 MCC 22.12 22.58 28.61 29.06 42.52 35.19 47.03 48.74 46.54 67.05 62.90 70.92 N2 Sleep Balanced Accuracy 64.85 67.47 65.88 67.72 Precision 73.86 74.95 76.25 76.33 Recall 75.49 76.27 77.22 77.50 F1 73.44 75.03 74.80 75.82 Cohen’s 33.51 37.98 36.71 39.58 MCC 35.28 38.87 39.68 41.29 N3 Sleep N3 Sleep Balanced Accuracy 76.48 78.15 79.40 81.16 77.36 81.50 80.08 82.18 Precision 89.80 90.53 90.86 91.71 90.46 90.98 90.92 91.72 Recall 85.36 87.47 87.60 89.66 88.07 81.92 86.85 88.59 F1 87.03 88.64 88.82 90.45 89.02 84.87 88.32 89.72 Cohen’s 39.73 45.03 46.36 52.35 45.54 38.84 45.47 50.70 MCC 41.84 46.52 48.08 53.44 46.58 44.30 47.77 52.60 REM Sleep REM Sleep REM Sleep Balanced Accuracy 75.29 81.55 82.31 86.60 77.83 77.66 82.99 88.67 70.09 85.57 84.13 89.26 Precision 90.95 93.38 92.70 94.69 91.67 93.31 92.63 94.97 90.33 93.47 92.52 94.87 Recall 90.33 93.24 91.39 94.31 90.93 93.76 90.70 94.26 91.04 91.80 89.04 93.81 F1 90.61 93.31 91.91 94.46 91.26 93.46 91.42 94.52 90.63 92.40 90.25 94.18 Cohen’s 47.59 61.83 56.59 69.06 51.59 60.63 55.18 70.06 43.76 60.09 51.95 68.78 MCC 47.72 61.85 57.27 69.23 51.80 61.06 56.35 70.54 44.03 61.23 54.38 69.61
Appendix G: Additional Model Performance Plots

### v4

@@ 4.3 Model Performance
The RNN model architecture obtained maximal per-class performances for wake and REM, and minimal per-class performances for N1 and N2 sleep. As a transitory state, periods of N1 sleep are short and instable, making its discrimination a difficult task. Physiological similarities between N2 and N3 sleep, such as cardiorespiratory signatures, similarly complicate the model’s ability to differentiate them, as evidenced by N2–N3 confusion as the most frequent misprediction for 5-class models (as well as N1/N2–N3 confusion for 4-class models). Of note, these classes also represent those hardest to distinguish from each other in sleep staging by human annotators [rosenberg:2013]. Conversely, wake and REM sleep have clearer differentiating characteristics, such as higher heart rate, which allows for easier classification.
No statistically significant variations in accuracy based on age, sex, PLMI, BMI, TST, sleep efficiency, insomnia, or RLS were exhibited by the 5-class ensemble model, providing an important advantage when applied in clinical contexts. The marginally significant () association between AHI and accuracy may owe to the fact that individuals with sleep apnea experience sleep fragmentation, resulting in N1 sleep, one of the most difficult stages to separate; nonetheless, the association was modest (). The model showed poorer performance in individuals with RBD, driven by difficulties in classifying REM sleep. This may reflect in part the centrality of accelerometry to our model, which, in the context of abnormal movements in REM sleep, may be apt to misclassify REM sleep.
The series of ablation analyses establish that much of the RNN model’s performance was likely fueled by chest module features (i.e. ECG, temperature, and XYZ accelerometry) rather than limb module (i.e. PPG and temperature), meaning that omitting the limb module may be reasonable if participant burden is a concern and collection of oximetry data is not necessary. Exclusion of both ECG and PPG data with retention of only accelerometry and temperature data resulted in model performance almost indistinguishable from use of all the chest module sensors, suggesting that the inclusion of traditional heart rate sensors (PPG, ECG) is not necessary. This highlights the richness of accelerometry data obtained at the chest, which may capture respiratory data as well as measures of movement, and may also contribute to the robustness of our model given the relative reliability of the accelerometers compared to the ECG/PPG sensors. At the same time, the attenuation of low-frequency activity (S.9) produced only modest degradations in performance, suggesting that the model is exploiting a nuanced combination of low-frequency respiratory motion, postural changes, and other movements, rather than any single narrow frequency band alone.
@@ 4.4 Limitations, Strengths, and Future Directions
The core strengths of this study are: 1) the ANNE One system itself, with multiple channels of data recorded from a pair of minimally intrusive sensors, is a strength insofar as it provides an element of redundancy should one sensor return low quality signal, 2) the inclusion of Mamba allowed the achievement of relatively high performance with a relatively parameter-scarce architecture, and 3) the study population was large and varied, including participants from across a spectrum of sleep disorders.
A few methodological limitations are also worth considering. 1) The black box nature of the RNN makes it difficult to ascertain the specific mechanisms used to infer sleep stage from the sensor data. 2) The recordings were overnight with a brief period of wakefulness before lights out, resulting in a preponderance of sleep vs. wake, so it is difficult to know for sure how the model will perform in detecting daytime naps during long ambulatory recordings, and further studies with ambulatory recordings in clinical populations will be informative. 3) The recordings were all obtained from a single site, and additional work is needed to assess the extent to which these results will be generalizable to ANNE recordings made in different settings on different patient populations.
The choice of 1-second windows created both advantages and disadvantages. It reduced the number of frequency bins per window, saving on model parameter count, and inflated the number of epochs by thirtyfold, which Mamba’s long-term dependencies could effectively capture. However, this setup complicated interpretability: while high-frequency content ( Hz) emerged in the model’s frequency bin features, low-frequency content ( Hz) manifested as changes across time windows, making their effects more difficult to capture with feature importance techniques.
An important future direction will be to better delineate how chest accelerometry contributes so strongly to model robustness, particularly in the absence of ECG and PPG. Our results suggest that some of this may be related to low-frequency, possibly respiratory, activity. However, the model appears to be making use of other information as well, and defining the physiological basis of this will be critical to better understanding the potential and limitations of chest accelerometry-based sleep staging.
@@ 5 Conclusion
This work introduces a Mamba-based RNN ensemble model to automatically perform 3-class, 4-class, and 5-class sleep staging using data from the ANNE One dual sensor using sleep stages inferred from concurrent PSG recordings. The model meets or surpasses performance metrics in related literature when applied to our ANNE One dataset, especially for wearable devices, and demonstrates robustness across a wide range of clinical characteristics. Moreover, the model’s capability was achieved without prescreening and excluding large numbers of patients or epochs for reasons of signal quality. Compared to PSGs, the ANNE One sensor is far less intrusive of a device and procedure, which is significant for expanding the accessibility of sleep measurement for older and remote-living adults at greater risk of sleep-related illness. Our model provides a scalable, accurate minimally-intrusive approach to the ambulatory assessment of sleep staging without the need for EEG.
Financial Disclosure Statement
The authors have no financial conflicts of interest to disclose.
@@
Non-Financial Disclosure Statement
The authors have no non-financial conflicts of interest to disclose.
Acknowledgments
Special thanks to Prof. Paul Kushner and Prof. Dylan Jones from the Department of Physics at the University of Toronto for offering various forms of support for this research during its infancy.
Data Availability
All data will be made available through the Canadian Federated Research Data Repository at https://www.frdr-dfdr.ca/repo/ upon publication. All code will be made available at https://github.com/OSHS2019/ upon publication.
S Supplementary Materials
S.1 Mamba Architecture
S.1.1 Selective SSM
At the core of the Mamba architecture is the state space model (SSM). For some single channel input , -dimensional latent state , single channel output , and continuous time , a SSM is governed by a system of first-order linear ordinary differential equations (ODE):
| where , , , are matrices that parameterize the SSM. However, in the context of sequence modeling, we require the SSM to work with discrete time steps . To this end we apply zero-order hold (ZOH) discretization to and in (S1) such that: | ||||
| where is the ZOH discretization time step, which governs the magnitude of each discrete update to the system’s trajectory. As such, we can reformulate the SSM from a system of ODEs to a recurrence: | ||||
| Note that Mamba’s SSM still learns the parameters and rather than the discretized and directly, where ZOH discretization is applied during forward pass. The Mamba architecture also imposes an additional structural constraint that is diagonal.
The SSM as formulated by (S5) and (S6) is linear time-invariant (LTI), which means that all parameters remain static and input-independent across time steps. This makes encoding context-awareness and adaptive behavior challenging. Mamba’s selective SSM breaks LTI, building input dependence into (S5) and (S6) by reparameterizing and . For a length input sequence : | ||||
| where is a parameterized linear projection from the input to .
Mamba’s selective SSM additionally reparameterizes to make the ZOH discretization time step adaptable to the input such that it functions similarly to RNN gating mechanisms, which is when the magnitude of each hidden state update is modulated as a function of the input: | ||||
where .
For a multi-channel input with channels , the selective SSM is applied channel-wise with a total of hidden states. Notice that by making the selective SSM not LTI, the Mamba architecture’s every hidden state update is dependent on the entire history of input such that it can capture arbitrarily long-range dependencies across the entire sequence similar to the Transformer architecture.
S.1.2 Mamba Block
A Mamba block additionally wraps the selective SSM (S.1.1) around linear projections, residual connections, and convolutions. The implementation of a Mamba block used in this study [mambapy] is as shown on Figure S1.
S.1.3 Bidirectional Mamba Block
Our bidirectional Mamba block consists of two Mamba blocks (S.1.2). One models the sequence in chronological order, while the other models the sequence in reverse-chronological order, as shown on Figure S2.
S.2 Summary Statistics by Groups
@@ Acknowledgments
S.3 Model Architecture
Below is the printed output of a -class RNN model (for ; for binary models, ) with exact hyperparameter values. The selective SSM parameters are contained within x_proj and dt_proj. Softmax probabilities found using the output of this RNN model are used to represent the hypnodensities of the regular RNN.
AnneNet(
(net): MambaNet(
(mamba_forward): Mamba(
(layers): ModuleList(
(0-2): 3 x ResidualBlock(
(mixer): MambaBlock(
(in_proj): Linear(in_features=10, out_features=40, bias=True)
(conv1d): Conv1d(20, 20, kernel_size=(4,), stride=(1,), padding=(3,), groups=20)
(x_proj): Linear(in_features=20, out_features=33, bias=False)
(dt_proj): Linear(in_features=1, out_features=20, bias=True)
(out_proj): Linear(in_features=20, out_features=10, bias=True)
(norm): RMSNorm()
(mamba_backward): Mamba(...same parameters as mamba_forward...)
(pre_mlp): Sequential(
(0): BatchNorm1d(140, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Linear(in_features=140, out_features=100, bias=True)
(2): LeakyReLU(negative_slope=0.01)
(3): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(4): Linear(in_features=100, out_features=100, bias=True)
(5): LeakyReLU(negative_slope=0.01)
(6): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(7): Linear(in_features=100, out_features=100, bias=True)
(8): LeakyReLU(negative_slope=0.01)
(9): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(10): Linear(in_features=100, out_features=10, bias=True)
(post_mlp): Sequential(
(0): BatchNorm1d(20, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Linear(in_features=20, out_features=100, bias=True)
(2): LeakyReLU(negative_slope=0.01)
(3): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(4): Linear(in_features=100, out_features=100, bias=True)
(5): LeakyReLU(negative_slope=0.01)
(6): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(7): Linear(in_features=100, out_features=100, bias=True)
(8): LeakyReLU(negative_slope=0.01)
(9): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(10): Linear(in_features=100, out_features=n, bias=True)
The suite of decision tree architectures below are grouped by base model. Each configuration (separated by semicolons) is a separate model. All omitted parameters are set to the default values in Scikit-learn or XGBoost.
Histogram Gradient Boosting: default parameters; max depth = 5; learning rate = 0.05; learning rate = 0.0333; learning rate = 0.01
Random Forest: ( = 3, max depth = 7); ( = 5, max depth = 5); ( = 10, max depth = 3); ( = 10, max depth = 3)
Extra Trees: ( = 10, max depth = 3)
AdaBoost: = 3; = 5
XGBoost: ( = 5, = 0.5); ( = 10, = 0.5)
The final 3-class, 4-class, and 5-class RNN ensemble models all chose a histogram gradient boosting classifier with learning rate . In general, gradient boosting classifiers use many decision trees where tree is trained on the erroneous predictions of , which gradually refines performance as increases. For multi-class classification, regression trees compute logits for each class before applying softmax:
As is miniscule, the RNN ensemble’s logits appear largely dominated by the initial prediction from the baseline decision tree . The first tree makes simple and underfitted predictions, resulting in stable but highly-uncertain probabilities such as those in Figure 3.
S.4 Additional Model Visualizations
UMAP projections on Figure S4 are generated using the following hyperparameters:
metric, output_metric: euclidean
n_neighbors: 15
n_components: 2
learning_rate: 1.0
min_dist: 0.1
spread: 1.0
set_op_mix_ratio: 1.0
local_connectivity: 1
negative_sample_rate: 5
target_n_neighbors: -1
target_weight: 0.5
dens_lambda: 2.0
dens_frac: 0.3
dens_var_shift: 0.1
S.5 Class-Wise Evaluation Metrics
@@
5-Class Sleep Staging 4-Class Sleep Staging 3-Class Sleep Staging Metric (%) Regular Ensemble Regular Ensemble Regular Ensemble Wake Wake Wake Balanced Accuracy 86.47 86.19 87.93 87.87 87.97 88.27 Precision 87.64 88.44 88.34 88.67 88.24 88.82 Recall 87.64 88.04 88.32 88.70 88.12 88.85 F1 87.54 87.81 88.33 88.66 88.16 88.83 Cohen’s 73.93 74.41 75.74 76.32 75.45 72.89 MCC 74.11 75.17 75.74 76.37 75.49 76.74 N1 Sleep N1/N2 Sleep Non-REM Sleep Balanced Accuracy 67.99 69.62 71.58 73.47 81.03 85.42 Precision 86.74 87.09 75.37 75.61 81.90 85.50 Recall 81.74 79.31 75.06 75.81 80.99 85.41 F1 83.77 82.25 74.00 75.36 80.87 85.41 Cohen’s 27.21 26.46 45.39 48.25 62.01 70.83 MCC 28.61 29.06 47.03 48.74 62.90 70.92 N2 Sleep Balanced Accuracy 65.88 67.72 Precision 76.25 76.33 Recall 77.22 77.50 F1 74.80 75.82 Cohen’s 36.71 39.58 MCC 39.68 41.29 N3 Sleep N3 Sleep Balanced Accuracy 79.40 81.16 80.08 82.18 Precision 90.86 91.71 90.92 91.72 Recall 87.60 89.66 86.85 88.59 F1 88.82 90.45 88.32 89.72 Cohen’s 46.36 52.35 45.47 50.70 MCC 48.08 53.44 47.77 52.60 REM Sleep REM Sleep REM Sleep Balanced Accuracy 82.31 86.60 82.99 88.67 84.13 89.26 Precision 92.70 94.69 92.63 94.97 92.52 94.87 Recall 91.39 94.31 90.70 94.26 89.04 93.81 F1 91.91 94.46 91.42 94.52 90.25 94.18 Cohen’s 56.59 69.06 55.18 70.06 51.95 68.78 MCC 57.27 69.23 56.35 70.54 54.38 69.61
S.6 CRNN Model Architecture
The CRNN model reorganizes the input data into 30-second windows mirroring the 30-second windows used in visual annotation of PSG data according to AASM criteria. As shown in Figure S5, the CRNN model comprises:
1)
A convolutional component, consisting of four 1D residual blocks (see S5, right) which continually extract features and downsize the window size axis, flattening it into an embedding vector per each 30-second window.
2)
A recurrent component, consisting of a bidirectional (same as RNN model, S.1) three-layer Mamba block that processes the embedding vectors sequentially, followed by an element-wise multi-layer perceptron (MLP) that diminishes the feature-size axis into logits for each class.
Below is the printed output of the layers of an -class CRNN model (for ; for binary models, ) with exact hyperparameter values.
CRNN(
(conv_blocks): ModuleList(
(0): ResidualBlock(
(net1): Sequential(
(0): BatchNorm1d(42, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Conv1d(42, 42, kernel_size=(7,), stride=(1,), padding=(3,))
(2): LeakyReLU(negative_slope=0.01)
(net2): Sequential(
(0): BatchNorm1d(42, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Conv1d(42, 64, kernel_size=(7,), stride=(2,), padding=(3,))
(2): LeakyReLU(negative_slope=0.01)
(mlp): Sequential(
(0): Linear(in_features=106, out_features=64, bias=True)
(1): LeakyReLU(negative_slope=0.01)
(2): Linear(in_features=64, out_features=64, bias=True)
(3): LeakyReLU(negative_slope=0.01)
(4): Linear(in_features=64, out_features=64, bias=True)
(5): LeakyReLU(negative_slope=0.01)
(pool): MaxPool1d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
(1): ResidualBlock(
(net1): Sequential(
(0): BatchNorm1d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Conv1d(64, 64, kernel_size=(7,), stride=(1,), padding=(3,))
(2): LeakyReLU(negative_slope=0.01)
(net2): Sequential(
(0): BatchNorm1d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Conv1d(64, 84, kernel_size=(7,), stride=(2,), padding=(3,))
(2): LeakyReLU(negative_slope=0.01)
(mlp): Sequential(
(0): Linear(in_features=148, out_features=84, bias=True)
(1): LeakyReLU(negative_slope=0.01)
(2): Linear(in_features=84, out_features=84, bias=True)
(3): LeakyReLU(negative_slope=0.01)
(4): Linear(in_features=84, out_features=84, bias=True)
(5): LeakyReLU(negative_slope=0.01)
(pool): MaxPool1d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
(2): ResidualBlock(
(net1): Sequential(
(0): BatchNorm1d(84, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Conv1d(84, 84, kernel_size=(7,), stride=(1,), padding=(3,))
(2): LeakyReLU(negative_slope=0.01)
(net2): Sequential(
(0): BatchNorm1d(84, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Conv1d(84, 104, kernel_size=(7,), stride=(2,), padding=(3,))
(2): LeakyReLU(negative_slope=0.01)
(mlp): Sequential(
(0): Linear(in_features=188, out_features=104, bias=True)
(1): LeakyReLU(negative_slope=0.01)
(2): Linear(in_features=104, out_features=104, bias=True)
(3): LeakyReLU(negative_slope=0.01)
(4): Linear(in_features=104, out_features=104, bias=True)
(5): LeakyReLU(negative_slope=0.01)
(pool): MaxPool1d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
(3): ResidualBlock(
(net1): Sequential(
(0): BatchNorm1d(104, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Conv1d(104, 104, kernel_size=(7,), stride=(1,), padding=(3,))
(2): LeakyReLU(negative_slope=0.01)
(net2): Sequential(
(0): BatchNorm1d(104, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Conv1d(104, 126, kernel_size=(7,), stride=(2,), padding=(3,))
(2): LeakyReLU(negative_slope=0.01)
(mlp): Sequential(
(0): Linear(in_features=230, out_features=126, bias=True)
(1): LeakyReLU(negative_slope=0.01)
(2): Linear(in_features=126, out_features=126, bias=True)
(3): LeakyReLU(negative_slope=0.01)
(4): Linear(in_features=126, out_features=126, bias=True)
(5): LeakyReLU(negative_slope=0.01)
(pool): MaxPool1d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
(rnn): MambaNet(
(mamba_forward): Mamba(
(layers): ModuleList(
(0-2): 3 x ResidualBlock(
(mixer): MambaBlock(
(in_proj): Linear(in_features=126, out_features=504, bias=True)
(conv1d): Conv1d(252, 252, kernel_size=(4,), stride=(1,), padding=(3,), groups=252)
(x_proj): Linear(in_features=252, out_features=40, bias=False)
(dt_proj): Linear(in_features=8, out_features=252, bias=True)
(out_proj): Linear(in_features=252, out_features=126, bias=True)
(norm): RMSNorm()
(mamba_backward): Mamba(...same parameters as mamba_forward...)
(mlp): Sequential(
(0): BatchNorm1d(252, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(1): Linear(in_features=252, out_features=100, bias=True)
(2): LeakyReLU(negative_slope=0.01)
(3): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(4): Linear(in_features=100, out_features=100, bias=True)
(5): LeakyReLU(negative_slope=0.01)
(6): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
(7): Linear(in_features=100, out_features=100, bias=True)
(rnn_activation): LeakyReLU(negative_slope=0.01)
(mlp): Sequential(
(0): Linear(in_features=226, out_features=64, bias=True)
(1): LeakyReLU(negative_slope=0.01)
(2): Linear(in_features=64, out_features=64, bias=True)
(3): LeakyReLU(negative_slope=0.01)
(4): Linear(in_features=64, out_features=16, bias=True)
(5): LeakyReLU(negative_slope=0.01)
(6): Linear(in_features=16, out_features=n, bias=True)
S.7 Additional Model Performance Plots
![[Uncaptioned image]](figures/unhealthy-subset.png)
@@
S.8 Accelerometry Signals
We aggregate our pre-processed accelerometry signals from the ANNE One chest module into Eucliean Norm Minus One (ENMO)1 and z-angle2. On Figures S12, S13, S14, S15, and S16, we select five representative full night recordings of healthy individuals from our study cohort and plot them against the ground truth hypnogram to show how accelerometry informs sleep stage transitions at the individual level. We also show power spectral density (PSD) heatmaps stratified by sleep stage for each recording.
S.9 Accelerometry Signal Ablation Study
Metric (%) Classes Feature Set Balanced Accuracy Precision Recall F1 Cohen’s MCC Accel. + Temp. 63.85 70.68 63.35 65.03 51.97 53.12 Accel. 64.60 71.15 64.18 65.64 52.97 54.12 5 Accel. (0.5 Hz High-Pass) 57.40 67.00 60.84 62.31 48.21 49.12 [1pt/1pt] Accel. + Temp. 73.03 74.46 71.52 72.10 58.70 59.30 Accel. 73.99 75.27 72.64 73.14 60.21 60.77 4 Accel. (0.5 Hz High-Pass) 67.23 72.00 68.98 69.55 54.98 55.65 [1pt/1pt] Accel. + Temp. 82.78 84.26 83.18 83.51 71.71 71.56 Accel. 83.08 84.82 83.76 84.08 72.47 72.61 3 Accel. (0.5 Hz High-Pass) 76.11 81.48 80.13 80.60 66.45 66.64
S.10 Feature Importance by Sensor